OpenStudy (anonymous):

slope-intercept form for the line that passes through (-8,1)and is perpendicular to the y=2x-17

OpenStudy (anonymous):

slope = -1 / 2 = -1/2 using y-y1 = m (x-x1) y - 1 = (-1/2) (x - (-8)) y - 1 = -1/2 (x + 8) y = (-1/2) x - 3

OpenStudy (anonymous):

thank you